OpenStudy (anonymous):

Solve for x in the equation 0.5x plus 78.2= 287

OpenStudy (agent47):

0.5x+78.2=287 0.5x = 287-78.2 x = (287-78.2)/0.5

OpenStudy (anonymous):

thank you